OpenStudy (anonymous):

Help please!!!!!!! Create your own equation written in standard form. 1. Determine any point that is a solution and will be on this line. Justify your answer algebraically.

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Do you remember the standard form of a line? :)

OpenStudy (anonymous):

ax + by = c isnt it?

zepdrix (zepdrix):

ok good. So just choose some numbers for a, b and c to complete the first step.

zepdrix (zepdrix):

Or the step prior to step1 i guess :P heh

OpenStudy (anonymous):

2x + 5y = 10

zepdrix (zepdrix):

Ok good. Let's come up with an ordered pair that we can use with this line. Whatchu got? :U

OpenStudy (anonymous):

I am confuse in this one. how do I find solution from this equation?

OpenStudy (anonymous):

I dont know how will I able to determine any point that is a solution and will be on this line. Justify your answer algebraically.

zepdrix (zepdrix):

Let's worry about that in a sec!! :) First come up with a number value for x and y that we can use.

zepdrix (zepdrix):

How about (x,y) = (3,4) ? k? :o

zepdrix (zepdrix):

What we want to do is, Plug 3 in for x, and 4 in for y, If our equation simplfies to 10 = 10, (the equation holds true), then (3,4) `is` a solution to the equation.

OpenStudy (anonymous):

okay I thought I have to plug zero

zepdrix (zepdrix):

\[\Large\bf\sf 2x + 5y = 10 \]You plug in any pair of numbers for x and y. And then state whether or not the equation holds true for that pair.

OpenStudy (anonymous):

okay thanks so much for help!!! :) you are too awesome!!!! thanks a lot for help!!!!!!

zepdrix (zepdrix):

np! understand how to finish it up? c:

OpenStudy (anonymous):

Well pretty much I think so, lol :)

zepdrix (zepdrix):

So is (3,4) a solution to the equation? :o

OpenStudy (anonymous):

yes I got it

OpenStudy (anonymous):

:)

OpenStudy (the_fizicx99):

^_^ you could also write it in y = mx + b and see where it crosses the line in the middle :# but that wouldn't be Ax + By = C Just seems easier than using random values :P

OpenStudy (anonymous):

lol :D I know but I was quite confused with solution. thanks anyways

OpenStudy (anonymous):

thanks so much again @zepdrix !!!!!! :)

zepdrix (zepdrix):

So the coordinate pair (3,4) is NOT a solution :o Just making sure you understand that. Ok ok you said you got it +_+ I'll go lol

OpenStudy (anonymous):

actually I think I am not. I understood now how to find a solution from this kind equation :( sorry

OpenStudy (anonymous):

how solution exactly like? if thats not solution

zepdrix (zepdrix):

Solution is a pair of numbers that gives you 10 = 10 for this equation. To find a solution, Choose an easy number to work with for one of your variables, like x=0. \[\Large\bf\sf 2x+5y=10\]Plugging in x=0,\[\Large\bf\sf 2\cdot0+5y=10\]\[\Large\bf\sf 5y=10\]Do you understand how to solve for y from this point?

OpenStudy (anonymous):

you mean, I have to plug zero to find solution of the equation like this one

zepdrix (zepdrix):

zero will be `part of your solution`. Your solution consists of 2 parts, x=something y=something else

zepdrix (zepdrix):

To find one solution we're letting, x=0 y=?

zepdrix (zepdrix):

We have to solve the equation for y.

OpenStudy (anonymous):

5y/5 = 10/5 y = 2

zepdrix (zepdrix):

Ok good. To check our work, we'll plug in x=0, y=2.\[\Large\bf\sf 2\cdot 0+5\cdot2\quad=\quad10\]\[\Large\bf\sf 10\quad=\quad 10 \qquad\color{green}{\checkmark}\]

zepdrix (zepdrix):

So we've determined that (0,2) is a solution to the equation.

OpenStudy (anonymous):

it is solution for just y. Right?

zepdrix (zepdrix):

No. it is a solution the equation. x=0, y=2.

OpenStudy (anonymous):

oh yeah, I am so sorry for giving you trouble and keeping you here quite long. :( I apologize for that.

zepdrix (zepdrix):

lol :3

OpenStudy (anonymous):

Honestly, but thank you so much again for coming back and explaining me! :)

OpenStudy (anonymous):

You are truly really great :) thanks for tolerating me and helping me a lot

OpenStudy (anonymous):

and you are really cool too :) I really enjoyed learning from you

zepdrix (zepdrix):

+_+